# Unbiased Estimation of Central Moments in Unbalanced Two- and Three-Level Models

### v1

@@ Abstract
Keywords: Multilevel (hierarchical) models; unbalanced panels; variances; third central moments; fourth central moments.
@@ 1 Introduction
This paper derives closed-form unbiased estimators of central moments in unbalanced two- and three-level models under two averaging schemes: group-level averaging (weighting clusters equally) and observation-level averaging (weighting units equally). In unbalanced panels, failing to account for unequal subgroup sizes can bias moment estimators for the latent components. Group-level averaging is natural when the sampling unit is the group, whereas observation-level averaging is natural when the sampling unit is the individual or when larger groups should receive proportionally greater weight.
@@ 2 Two-level model
where $\{u_{i}\}_{i=1}^{n}$ and $\{v_{ij}\}_{i,j}$ are unobserved, mutually independent, and within each level i.i.d., with $\mathbb{E}[u]=\mathbb{E}[v]=0$. We assume $n\geq 3$ and $J_{i}\geq 3$ for all $i$. Let $\mu_{ku}=\mathbb{E}[u^{k}]$ and $\mu_{kv}=\mathbb{E}[v^{k}]$ for $k=2,3,4$, and define sample means (superscripts $\mathrm{g}$ and $\mathrm{o}$ denote group-level and observation-level averaging) as follows,
@@ 3 Three-level model
where $\{u_{i}\}_{i=1}^{n}$, $\{v_{ij}\}_{i,j}$, and $\{w_{ijk}\}_{i,j,k}$ are unobserved, mutually independent, and within each level i.i.d., with $\mathbb{E}[u]=\mathbb{E}[v]=\mathbb{E}[w]=0$. We assume $n\geq 3$, $J_{i}\geq 3$, and $K_{ij}\geq 3$ for all $(i,j)$. Let $\mu_{ku}=\mathbb{E}[u^{k}]$, $\mu_{kv}=\mathbb{E}[v^{k}]$, and $\mu_{kw}=\mathbb{E}[w^{k}]$ for $k=2,3$, and define sample means as follows,

### v2

@@ Abstract
Keywords: Multilevel (hierarchical) models; unbalanced panels; central moments; group-level averaging; observation-level averaging.
@@ 1 Introduction
This paper derives closed-form unbiased estimators of central moments in unbalanced two- and three-level models under two averaging schemes: group-level averaging (weighting groups equally) and observation-level averaging (weighting lowest-level units equally). In unbalanced panels, failing to account for unequal subgroup sizes can bias moment estimators for the latent components. The appropriate weighting depends on the object of interest: group-level averaging is natural when each group is meant to count equally, while observation-level averaging is natural when each lowest-level unit is meant to count equally.
@@ 2 Two-level model
where $\{u_{i}\}_{i}$ and $\{v_{ij}\}_{i,j}$ are unobserved, mutually independent, and within each level i.i.d., with $\mathbb{E}[u]=\mathbb{E}[v]=0$. We assume $n\geq 3$ and $J_{i}\geq 3$ for all $i$. Let $\mu_{ku}=\mathbb{E}[u^{k}]$ and $\mu_{kv}=\mathbb{E}[v^{k}]$ for $k=2,3,4$, and define sample means (superscripts $\mathrm{g}$ and $\mathrm{o}$ denote group-level and observation-level averaging) as follows,
@@
Under balanced designs ($J_{i}=J$ for all $i$), the group- and observation-level averaging estimators coincide. Under unbalanced designs, the difference between them is weighting: group-level averaging gives each group effect $u_{i}$ equal weight, while observation-level averaging weights group effects in proportion to $J_{i}$. Thus, the latter gives more influence to larger groups. This provides the main source of guidance to researchers: group-level averaging should be used when each group is meant to count equally, whereas observation-level averaging should be used when each lowest-level unit is meant to count equally. For example, if $u_{i}$ is a teacher effect and the goal is to describe the distribution of teacher quality across teachers, group-level averaging is natural. If instead the goal is to describe the distribution of teacher quality as experienced by a randomly sampled student, observation-level averaging is natural, since students are more likely to be assigned to larger classes. Group-level averaging may also be preferable when comparing moments across samples or periods, since otherwise changes in group sizes mechanically change the weights attached to the same group effects.
A second consideration is the variance of the estimator. To see the intuition, consider the second-moment estimators. Both use between-group variation in the sample means $\bar{y}_{i}$, but these means are measured with different precision: writing $\bar{y}_{i}=u_{i}+\bar{v}_{i}$, we have $\text{Var}(\bar{y}_{i})=\mu_{2u}+\mu_{2v}/J_{i}$, so smaller groups yield noisier sample means. Observation-level averaging therefore places more weight on larger groups, for which $\mu_{2v}/J_{i}$ is smaller. This suggests that, when group sizes are very unequal and $\mu_{2v}$ is large relative to $\mu_{2u}$, observation-level averaging may have smaller estimator variance. When $\mu_{2v}$ is small relative to $\mu_{2u}$, however, this variance advantage is attenuated. Because the two averaging schemes correspond to different objects of interest under unbalancedness, estimator variance is secondary to the choice of weighting.
@@ 3 Three-level model
where $\{u_{i}\}_{i}$, $\{v_{ij}\}_{i,j}$, and $\{w_{ijk}\}_{i,j,k}$ are unobserved, mutually independent, and within each level i.i.d., with $\mathbb{E}[u]=\mathbb{E}[v]=\mathbb{E}[w]=0$. We assume $n\geq 3$, $J_{i}\geq 3$, and $K_{ij}\geq 3$ for all $(i,j)$. Let $\mu_{ku}=\mathbb{E}[u^{k}]$, $\mu_{kv}=\mathbb{E}[v^{k}]$, and $\mu_{kw}=\mathbb{E}[w^{k}]$ for $k=2,3$, and define sample means as follows,